2^1465 ,3^1172,5^879 which is greater what is the greatest number

Grade:8

1 Answers

Arun
5 years ago

Answer is 5^879

This is how it can be easily solved

5^879, 3^1172, 2^1758,

= 5^(3*293), 3^(4*293), 2^(5*293),

= (5^3)^293, (3^4)^293, (2^5)^293,

= 125^293, 81^293, 32^293,

Now its simple...

5^879 is the greatest :)